RC meeting and Run Plan, Thursday Oct. 5
- We got 43% of min stat. goal at Q2=2.115 W=2.95 x=0.21 as planned
- Changed SHMS aerogel tray ; ready for beam permit by ~ 1 pm
- Howard fixed ringing of wires in HMS DC 
- Steve L., Jack and Ethan fixed the issue with the HMS Dipole LN2 JT valve; no need to regulate manually
- Beam current not high/stable enough
- Plans - Physics '
- Take data at Q2=4.4 W=2.74 x=0.4: theta_Shms=9.81, theta_Shms=12.81, theta_Shms=15.81 (roughly 2 shifts per SHMS angle)